# Humidity-Aware Model Predictive Control for Residential Air Conditioning: A Field Study

### v1

@@ Abstract
Model predictive control of residential air conditioning could reduce energy costs and greenhouse gas emissions while maintaining or improving occupants’ thermal comfort. However, most approaches to predictive air conditioning control either do not model indoor humidity or treat it as constant. This simplification stems from challenges with modeling indoor humidity dynamics, particularly the high-order, nonlinear equations that govern heat and mass transfer between the air conditioner’s evaporator coil and the indoor air. This paper develops a machine-learning approach to modeling indoor humidity dynamics that is suitable for real-world deployment at scale. This study then investigates the value of humidity modeling in four field tests of predictive control in an occupied house. The four field tests evaluate two different building models: One with constant humidity and one with time-varying humidity. Each modeling approach is tested in two different predictive controllers: One that focuses on reducing energy costs and one that focuses on constraining electric power during a demand response event. The two models lead to similar performance for reducing energy costs. Combining the results of this study and a prior heating study of the same house, the estimated year-round energy cost savings were \$340–497 or 22–31% (95% confidence intervals); these savings were consistent across both humidity models. However, in the demand response tests, the simplifying assumption of constant humidity led to far more frequent and severe violations of the power constraint. These results suggest that accurate building models are important for nonlinear objectives, such as reducing or constraining peak demand, while for linear objectives such as reducing energy costs or emissions, model accuracy is less important.
@@ 1.1 Supervisory air conditioning control
Supervisory control of heating, ventilation, and air conditioning (HVAC) equipment is one way to improve the energy efficiency of air conditioning [5, 6]. Supervisory HVAC control systems dynamically adjust set-points – such as indoor air temperatures, compressor speeds, or fan speeds – that device-level control systems track. One prevalent supervisory control methodology is model predictive control (MPC) [7, 8]. This approach, illustrated in Fig. 1, typically uses models of the HVAC equipment and building, as well as forecasts of weather and occupancy, to decide control actions that optimize performance objectives over a receding prediction horizon [5].
@@
Air conditioners affect, and are affected by, both the temperature and humidity of the indoor air. Pergantis et al. [9] recently reviewed the literature on field demonstrations of supervisory HVAC control in residential buildings and found that no air conditioning study accounted for time-varying indoor humidity and dehumidification load. However, in commercial buildings, other researchers have shown in numerous simulation studies [10, 11], as well as three experimental studies [12, 13, 14], that neglecting humidity can decrease comfort and controller performance. These studies modeled the indoor humidity as a separate state in a hygrothermal model, impacted by the indoor dry-bulb temperature and HVAC dehumidification. However, there is significant engineering effort in training the multiple correlations to model the indoor heat exchanger [10, 15], as well as in solving the resulting nonconvex optimization problem [10]. Several review papers [16, 17, 18] have noted the challenges related to modeling dehumidification in HVAC systems, as well as the lack of humidity modeling in many simulation studies [19, 20, 21]. For these reasons, most experimental demonstrations of supervisory air conditioning control model only temperature dynamics and make the simplifying assumption that humidity is constant [22, 23, 24, 25].
@@
This paper presents the first experimental evaluation of an MPC-based supervisory controller in a residential building that considers dehumidification in the problem formulation. A simplified approach is developed to reduce the complexity of training a humidity model and modeling the indoor heat exchanger. In this approach, the indoor humidity state, defined here as the return air (mixed from the whole building) wet-bulb temperature, is predicted through machine learning. This approach is well suited for residential buildings and low-cost supervisory controllers. Two methods are formulated to predict an air conditioner’s total electricity demand, which can be decomposed into latent (dehumidifaction) and sensible (temperature-change) components. The first method generates load predictions wherein the sensible heat ratio (SHR, the ratio of sensible load to total load) varies over time. The second method, by contrast, assumes a constant SHR. Following the naming convention of Raman et al. [10, 15], we refer to the first load prediction method as the “latent” model and the second as the “sensible” model. Each load prediction model is tested in two MPC implementations. The first MPC implementation aims to minimize energy cost and maximize thermal comfort. The second MPC implementation also aims to minimize energy and maximize comfort, but also seeks to constrain electrical power below a threshold between 4 PM and 8 PM. This power constraint emulates a demand response call during the hours when residential electricity demand typically peaks.
@@ 2.1 Simulations
Heat exchanger modeling. One of the major challenges of modeling humidity dynamics in buildings arises from the complex dehumidification phenomena occurring across the indoor heat exchanger (evaporator). One class of models involves solving complex partial differential equations based on the geometry of the heat exchanger and its different regions (e.g., dry, wet, partial conditions) [36]. Since time constants are typically on the order of hours for building thermal dynamics and on the order of minutes for the heat exchanger [10, 36], static heat exchanger models suffice. To extend the capabilities of static models as the inlet conditions change, Raman et al. [10] used 1,159 temperature and humidity binds to characterize the inlet conditions to the heat exchanger, with each bin using a 5th-degree polynomial model. The resulting root-mean-square error (RMSE) for the prediction of the supply air temperature and relative humidity were 0.28 C and 1%, respectively. The model was trained using EnergyPlus [37]. Other studies have similarly constructed multiple regression models based on the temperature and humidity operating regions [38] or have simplified governing equations into large sparse systems of linear ODEs [39].
@@ 3.1 Building modeling
A derivation of the discrete-time governing equations and the model training process can be found in [9]. The model was trained the model on passive observations from June 20 to July 20, 2023. These 30 days gave 720 hourly data points for each observed variable. One-quarter of the training set included days under MPC operation; the other three-quarters of the training data came from the manufacturer’s default control algorith with constant indoor temperature set-points. Fig. 4 shows the end-to-end fit, including the thermal circuit and the exogenous thermal power predictor, in training and validation data. The one-step-ahead indoor temperature predictions (left plot) match the targets with RMSE of 0.53 C in the validation data. The sensible cooling rate predictions matched the measurements with an RMSE of 3.6 kW. Over a steady thermostat set-point week, the RMSE for the power prediction is 0.44 kW. Overall, the accuracy of the temperature prediction was found to be sufficient for control purposes, consistent with findings in [47].
@@ 3.2 Equipment and humidity models
This study uses two different formulations for the heat pump coefficient of performance (COP), both based on manufacturer data. The formulations are part of a supervisory control system (MPC) that adjusts indoor temperature set-points. In the “sensible” formulation, the SHR is assumed constant at 0.86. Under this condition, the COP is modeled only as a function of the outdoor air temperature. After the completion of MPC testing, the actual mean SHR was found to be 0.79 over the testing duration. The impact of this assumption is considered in the discussion section. In the “latent” model formulation, the SHR varies during the day. By reviewing the manufacturer’s data, it was found that the SHR can be approximated as a linear function of the heat exchanger inlet air wet-bulb temperature. Additionally, the COP in the “latent” formulation is considered to be a function of both the outdoor air dry-bulb temperature and the indoor air wet-bulb temperature, as shown in Fig. 5.
Figure 5: A quadratic model of the HVAC coefficient of performance as a function of indoor wet-bulb and outdoor dry-bulb temperatures fits the training data with $R^{2}$ = 0.99.
@@
$T_{\text{wb}}=\text{GPR}(RH_{\text{out}},T_{out},h,I_{\text{solar}},w)$ (1)
@@
with the constants $a$ and $b$ fit using linear regression. Using the forecasted wet-bulb temperature, the SHR is obtained with an RMSE of 0.1 ($R^{2}$ = 0.95). The performance of the model in a typical week is shown in Fig. 7.
Figure 6: A linear model of SHR vs. indoor wet-bulb temperature fits the training data with $R^{2}$ = 0.97. Variations are due to different outdoor temperatures.
@@ 3.3 Supervisory controller design
Optimization inputs. In field demonstrations of the supervisory control system, a time step duration of $\Delta t=1$ h was used, prediction horizon of $L=24$ time steps, constant electrical energy price $\pi_{e}=0.14$/kWh, heat pump capacity $P_{\text{HP,max}}=4.5$ kW, temperature deviation $\delta=3$ C, and trained discrete-time dynamics parameter $\alpha=0.86$. The thermal resistance was $R=1.04$ C/kW, with the outdoor conditions prediction obtained from weather forecasts. The COP is predicted in the case of the “sensible” formulation by propagating the outdoor temperature forecast through a quadratic function fit to manufacturer data, while the “latent” formulation uses both the outdoor temperature forecast and the GPR prediction of the indoor wet-bulb temperature. The exogenous thermal power $\dot{Q}_{e}$ is predicted by propagating weather predictions and time features through the support vector machine described in [9]. The SHR is assumed to be a constant value of 0.86 for the “sensible” model formulation and is predicted in the “latent” model formulation using the linear model in Fig. 6. The occupants specified preference temperatures $T_{\text{pref}}$ of 23 C, constant during the day. A daytime setback was not considered since the site has frequent visitors. A flat electricity price was used since that is the utility structure in place in West Lafayette, Indiana, where the test house resides. However, the control framework can accommodate any dynamic pricing scheme, such as time-of-use or real-time pricing. The optimization problem presented in Eq. (3) is convex and solved in Matlab using the CVX toolbox [48]. Infeasibilities rarely occur due to the soft treatment of the temperature bounds, but when they do, the control system defaults to the indoor temperature set-point from the previous time step.
Price tuning. The formulation presented in Eq. (3a) involves prices that govern trade-offs between multiple objectives, such as peak electricity demand, energy costs, and thermal comfort. Achieving good controller performance requires tuning the peak demand price $\pi_{d}$ and thermal discomfort price $\pi_{t}$. The utility serving the test-site does not impose a peak demand charge. Rather, the demand price is included as an additional incentive to mitigate demand peaks. Tuning the peak demand price $\pi_{d}$ to \$0.8 per kW of daily peak demand, corresponding to a monthly peak demand price of about 25 \$/kW (a typical value for monthly peak demand charges in commercial buildings in the USA) gave good performance.
@@ 3.4 Thermal circuit model re-calibration
Upon the completion of the cost-reducing MPC testing (24 days), it was found that although the controller can provide energy savings, the building model (Fig. 4) was unable to capture the true electrical peaks. This is shown across a typical week of testing in Fig. 9 where the model predictions are compared to the real values from the site. This was attributed to overfitting parameters, particularly the effective thermal resistance, to the training data. However, it was found that even with this limitation, the controller was able to achieve energy savings (Sections 5.3, 5.4, and Fig. 11). This corroborates the findings of other researchers [47, 49] that as long the controller can reasonably predict the temperature at the next time-step and the trajectory of the load, predicting the magnitude of the load with high accuracy is not necessary. This does not apply in cases where objectives require accurate model predictions, such as minimizing peak electrical power usage or maintaining power below a specific target. The right plot in Fig. 9 shows an improved thermal circuit fit, which involved manual tuning of the physical parameters. For the improved fit, $\alpha=0.77$, $R=0.42$ C/kW, and $\dot{Q}_{e}$ = 3.4 kW (constant) were used. The resulting model has an improved prediction of the one-step-ahead temperature (RMSE of 0.4 C), and while the prediction of the cooling rate is slightly worse (RMSE of 4.2 kW), it does not attempt to fit the mean of the data as the previous model. It can be seen that with a poor prediction of the sensible load (Fig. 9, left), the differences in the two model formulations are not discernable, while with the improved thermal circuit model (Fig. 9, right), the “latent” model formulation better captures the true electrical peaks.
@@ 4.1 Test House
This section briefly introduces the DC Nanogrid House (Fig. 1) as an experimental testbed; Pergantis et al. [9] described the house and instrumentation in more detail. The DC Nanogrid House is a 208 m2, two-story, 1920s-era detached single-family home near Purdue University’s campus in West Lafayette, USA. This location falls under the International Energy Conservation Climate Code 5A. This climate sees both hot, humid summers (up to 35 C) and cold winters (down to $-20$ C). The DC Nanogrid House’s exterior walls have foam insulation with an R-Value of 3.5 C m2/W. Code-minimum U-8 W/m2/C windows make up about 20% of the exterior wall area. Networked sensors throughout the thermal and electrical systems have been installed. The house is a living laboratory occupied by graduate students.

### v2

@@ Abstract
Model predictive control of residential air conditioning could reduce energy costs and greenhouse gas emissions while maintaining or improving occupants’ thermal comfort. However, most approaches to predictive air conditioning control either do not model indoor humidity or treat it as constant. This simplification stems from challenges with modeling indoor humidity dynamics, particularly the high-order, nonlinear equations that govern heat and mass transfer between the air conditioner’s evaporator coil and the indoor air. This paper develops a machine-learning approach to modeling indoor humidity dynamics that is suitable for real-world deployment at scale. This study then investigates the value of humidity modeling in four field tests of predictive control in an occupied house. The four field tests evaluate two different building models: One with constant humidity and one with time-varying humidity. Each modeling approach is tested in two different predictive controllers: One that focuses on reducing energy costs and one that focuses on constraining electric power below a utility-specified threshold. The two models lead to similar performance for reducing energy costs. Combining the results of this study and a prior heating study of the same house, the estimated year-round energy cost savings were \$340–497 or 22–31% (95% confidence intervals); these savings were consistent across both humidity models. However, in the demand response tests, the simplifying assumption of constant humidity led to far more frequent and severe violations of the power constraint.
@@ 1.1 Supervisory air conditioning control
Supervisory control of heating, ventilation, and air conditioning (HVAC) equipment is one way to improve the energy efficiency of air conditioning [5, 6]. Supervisory HVAC control systems dynamically adjust set-points – such as indoor air temperatures, compressor speeds, or fan speeds – that device-level control systems track. One prevalent supervisory control methodology is model predictive control (MPC) [7, 8]. This approach, illustrated in Fig. 1, typically uses models of the HVAC equipment and building, as well as forecasts of weather and occupancy, to repeatedly decide control actions that optimize performance objectives over a receding prediction horizon [5].
@@
Air conditioners affect, and are affected by, both the temperature and humidity of the indoor air. The extent to which humidity modeling influences controller performance is an open question. The answer likely depends on the equipment configuration, building type, climate zone, and control objectives. Pergantis et al. [9] recently reviewed the literature on field demonstrations of supervisory HVAC control in residential buildings and found that no air conditioning study accounted for time-varying indoor humidity and dehumidification load. However, in commercial buildings, other researchers have shown in numerous simulation studies [10, 11], as well as three experimental studies [12, 13, 14], that neglecting humidity can decrease comfort and controller performance. These studies modeled the indoor humidity as a separate state in a hygrothermal model, impacted by the indoor dry-bulb temperature and HVAC dehumidification. However, there is significant engineering effort in training the multiple correlations to model the indoor heat exchanger [10, 15], as well as in solving the resulting nonconvex optimization problem [10]. Several review papers [16, 17, 18] have noted the challenges related to modeling dehumidification in HVAC systems, as well as the lack of humidity modeling in many simulation studies [19, 20, 21]. For these reasons, most experimental demonstrations of supervisory air conditioning control model only temperature dynamics and make the simplifying assumption that humidity is constant [22, 23, 24, 25], neglecting humidity dynamics driven by time-varying occupant behavior.
@@
This paper presents the first experimental evaluation of an MPC-based supervisory controller in a residential building that considers dehumidification in the problem formulation. A simplified approach is developed to reduce the complexity of training a humidity model and modeling the indoor heat exchanger. In this approach, the indoor humidity state, defined here as the return air (mixed from the whole building) wet-bulb temperature, is predicted through machine learning. This approach is well-suited for residential buildings, where active humidity control is typically not required [26], but where better predicting the impact of humidity on energy use can improve controller performance. Two methods are formulated to predict an air conditioner’s total electricity demand, which can be decomposed into latent (humidity-change) and sensible (temperature-change) components. The first method generates load predictions wherein the sensible heat ratio (SHR, the ratio of sensible load to total load) varies over time. The second method, by contrast, assumes a constant SHR. Following the naming convention of Raman et al. [10, 15], we refer to the first load prediction method as the “latent” model and the second as the “sensible” model. Each load prediction model is tested in two MPC implementations, resulting in a total of four experiments. The first MPC implementation aims to minimize energy cost and maximize thermal comfort. The second MPC implementation also aims to minimize energy and maximize comfort, but also seeks to constrain electrical power below a utility-specified threshold between 4 PM and 8 PM. This power constraint emulates a demand response call during the hours when residential electricity demand typically peaks. For energy cost minimization, the latent and sensible models give similar performance, which corroborates the findings of previous authors [27, 28] on the impact of model accuracy on energy savings. For power limiting, although both latent and sensible models enable some load shifting, the sensible model leads to significantly more frequent, longer, and more severe constraint violations.
@@ 2.1 Simulations
Heat exchanger modeling. One of the major challenges of modeling humidity dynamics in buildings arises from the complex dehumidification phenomena occurring across the indoor heat exchanger. One class of models involves solving systems of partial differential equations based on the geometry of the heat exchanger and its dry, wet, and mixed regions [39]. The structure and parameters of the governing equations vary between types of heat exchangers, such as air-to-water or air-to-refrigerant [39].
Since time constants are typically on the order of hours for building thermal dynamics and on the order of minutes for the heat exchanger [10, 39], static heat exchanger models suffice. To extend the capabilities of static models as the inlet conditions change, Raman et al. [10] used 1,159 temperature and humidity binds to characterize the inlet conditions to the heat exchanger, with each bin using a 5th-degree polynomial model. The resulting root-mean-square error (RMSE) for the prediction of the supply air temperature and relative humidity were 0.28 C and 1%, respectively. The model was trained using EnergyPlus [40]. Other studies have similarly constructed multiple regression models based on the temperature and humidity operating regions [41] or have simplified governing equations into large sparse systems of linear ODEs [42].
@@ 2.3 Research gaps
Figure 3: The workflow used in this paper to model and control the building and equipment.
@@ 3 Modeling, learning, and control methods
This section presents the methods used for modeling the system dynamics, training machine-learning models, and controlling the HVAC equipment. Section 3.1 describes the process of identifying a low-order resistance-capacitance model of a detached house with a single mixed-air zone. Section 3.2 develops two humidity models: (1) a latent model that considers a time-varying indoor wet-bulb temperature for computing the air conditioner’s coefficient and SHR, and (2) a sensible model that assumes a constant SHR and models the coefficient of performance (COP) as a function only of the outdoor temperature. Finally, Section 3.3 develops two separate supervisory controller formulations: one aimed at reducing energy costs, and one aimed at constraining power below a utility-specified threshold during a demand response window. Fig. 3 illustrates the workflow of model development and control design presented in this section.
@@ 3.1 Building modeling
A derivation of the discrete-time governing equations and the model training process can be found in [9]. The model was trained the model on passive observations from June 20 to July 20, 2023. These 30 days gave 720 hourly data points for each observed variable. One-quarter of the training set included days under MPC operation; the other three-quarters of the training data came from the manufacturer’s default control algorith with constant indoor temperature set-points. Fig. 5 shows the end-to-end fit, including the thermal circuit and the exogenous thermal power predictor, in training and validation data. The one-step-ahead indoor temperature predictions (left plot) match the targets with RMSE of 0.53 C in the validation data. The sensible cooling rate predictions matched the measurements with an RMSE of 3.6 kW. Over a steady thermostat set-point week, the RMSE for the power prediction is 0.44 kW. The model occasionally predicts heat demand on the order of 700-1000 W. While these heating predictions are unrealistic during the cooling season, the accuracy of the temperature prediction was found to be sufficient for control purposes. The observed suitability of a moderately accurate model for predictive control is consistent with prior findings, such as [27].
@@ 3.2 Equipment and humidity models
This study uses two different formulations for the heat pump COP, both based on manufacturer data. The formulations are part of a supervisory control system (MPC) that adjusts indoor temperature set-points. In the sensible formulation, the SHR is assumed constant at 0.86, consistend with the heat pump manufacturer’s specification sheet. Under this condition, the COP is modeled only as a function of the outdoor air temperature. After the completion of MPC testing, the actual mean SHR was found to be 0.79 over the testing duration. The impact of this assumption is considered in the discussion section. In the latent model formulation, the SHR varies during the day. By reviewing the manufacturer’s data, it was found that the SHR can be approximated as a linear function of the heat exchanger inlet air wet-bulb temperature. Additionally, the COP in the latent formulation is considered to be a function of both the outdoor air dry-bulb temperature and the indoor air wet-bulb temperature, as shown in Fig. 6.
Figure 6: A quadratic model of the HVAC COP as a function of indoor wet-bulb and outdoor dry-bulb temperatures fits the training data with $R^{2}$ = 0.99. Performance data were considered at an indoor dry-bulb temperature of 24 C, matching historical set-points on-site.
@@
Figure 7: A linear model of SHR vs. indoor wet-bulb temperature fits the training data with $R^{2}$ = 0.97. Variations are due to different outdoor dry-bulb temperatures.
@@
$T_{\text{wb}}=\text{GPR}(RH_{\text{out}},T_{out},h,I_{\text{solar}},w).$ (1)
@@
with the constants $a$ and $b$ fit using linear regression. The training data provided by the manufacturer included the sensible and total cooling rates at a fixed indoor dry-bulb temperature under varying indoor wet-bulb temperatures and outdoor dry-bulb temperatures. The forecasted wet-bulb temperatures and the linear heat exchanger regression in Eq. (2) gave a SHR prediction RMSE of 0.1 and $R^{2}$ of 0.95. Fig. 8 shows the performance of the model in a typical week.
@@ 3.3 Supervisory controller design
Optimization inputs. In field demonstrations of the supervisory control system, a time step duration of $\Delta t=1$ h was used, prediction horizon of $L=24$ time steps, constant electrical energy price $\pi_{e}=0.14$/kWh, heat pump capacity $P_{\text{HP,max}}=4.5$ kW, temperature deviation $\delta=3$ C, and trained discrete-time dynamics parameter $\alpha=0.86$. The temperature deviation band was calibrated in simulation to ensure acceptable indoor conditions and PPD across varying ambient temperatures and humidities. The thermal resistance was $R=1.04$ C/kW, with the outdoor conditions prediction obtained from weather forecasts. The COP is predicted in the case of the sensible formulation by propagating the outdoor temperature forecast through a quadratic function fit to manufacturer data, while the latent formulation uses both the outdoor temperature forecast and the GPR prediction of the indoor wet-bulb temperature. The exogenous thermal power $\dot{Q}_{e}$ is predicted by propagating weather predictions and time features through the support vector machine described in [9]. The SHR is assumed to be a constant value of 0.86 for all time steps for the sensible model formulation and is predicted in the latent model formulation using the linear model in Fig. 7. The occupants specified preference temperatures $T_{\text{pref}}$ of 23 C, constant during the day. A daytime setback was not considered since the site has frequent visitors. A flat electricity price was used since that is the utility structure in place in West Lafayette, Indiana, where the test house resides. However, the control framework can accommodate any dynamic pricing scheme, such as time-of-use or real-time pricing. The optimization problem presented in Eq. (3) is convex and solved in Matlab using the CVX toolbox [50]. Infeasibilities rarely occur due to the soft treatment of the temperature bounds, but when they do, the control system defaults to the indoor temperature set-point from the previous time step.
Price tuning. The formulation presented in Eq. (3a) involves prices that govern trade-offs between multiple objectives, such as peak electricity demand, energy costs, and thermal comfort. Achieving good controller performance requires tuning the peak demand price $\pi_{d}$ and thermal discomfort price $\pi_{t}$. The utility serving the test-site does not impose a peak demand charge. Rather, the demand price is included as an additional incentive to mitigate demand peaks or provide power-limiting control. Tuning the peak demand price $\pi_{d}$ to \$0.8 per kW of daily peak demand, corresponding to a monthly peak demand price of about 25 \$/kW (a typical value for monthly peak demand charges in commercial buildings in the USA) gave good performance.
@@ 3.4 Thermal circuit model re-calibration
Upon the completion of the cost-reducing MPC testing (24 days), it was found that although the controller can provide energy savings, the building model (Fig. 5) was unable to capture the true electrical peaks. This is shown across a typical week of testing in Fig. 10 where the model predictions are compared to the real values from the site. However, despite this, the controller was able to achieve energy savings (Sections 5.3, 5.4, and Fig. 12). This corroborates the findings of other researchers [27, 28] that as long the controller can reasonably predict the temperature at the next time step and the trajectory of the load, predicting the magnitude of the load with high accuracy is not necessary. This may not apply in cases that require very accurate power predictions, such as minimizing peak demand, constraining power below a specific target, or controlling thermally massive systems [51].
The right plot in Fig. 10 shows an improved thermal circuit fit, which involved manual tuning of the physical parameters. For the improved fit, $\alpha=0.77$, $R=0.42$ C/kW, and $\dot{Q}_{e}$ = 3.4 kW (constant) were used. The resulting model has an improved prediction of the one-step-ahead temperature (RMSE of 0.4 C), and while the prediction of the cooling rate is slightly worse (RMSE of 4.2 kW), it does not attempt to fit the mean of the data as the previous model.
@@ 4.1 Test House
This section briefly introduces the DC Nanogrid House (Fig. 1) as an experimental testbed; Pergantis et al. [9] described the house and instrumentation in more detail. The DC Nanogrid House is a 208 m2, two-story, 1920s-era detached single-family home near Purdue University’s campus in West Lafayette, USA. This location falls under the International Energy Conservation Climate Code Zone 5A. This climate sees both hot summers (up to 35 C) and cold winters (down to $-20$ C). Climate Zone 5A is also humid. At the test site, for example, over 25% of the experiment days had a dewpoint temperature greating than 20 C, indicating humid conditions [52].
The DC Nanogrid House’s exterior walls have foam insulation with an R-Value of 3.5 C m2/W. Code-minimum U-8 W/m2/C windows make up about 20% of the exterior wall area. Networked sensors throughout the thermal and electrical systems have been installed. The house is a living laboratory occupied by graduate students.